# Instability thresholds for de Sitter and Minkowski spacetimes in holographic semiclassical gravity

### v1

@@ 4 The perturbed solutions with negative mass-squared
$\delta\langle~{{\cal T}_{\mu}}^{\nu}$$\rangle$ $=$ $-\frac{\epsilon L^{4}p}{144\pi G_{6}\ell^{5}}p\,\hat{m}^{2}(\hat{m}^{2}-3){H_{\mu}}^{\nu}$ (4.20)
@@ 5.1.1 Stability analysis in the static region of de Sitter spacetime
$ds_{d}^{2}$ $=$ $-\frac{u-1}{u}(1+\epsilon T(u)e^{-i\omega t})dt^{2}+\frac{\epsilon\ell S(u)e^{-i\omega t}}{2u^{2}(u-1)}dtdu$ (5.1)
$+\frac{\ell^{2}}{4u^{2}(u-1)}(1+\epsilon U(u)e^{-i\omega t})du^{2}+\frac{\ell^{2}}{u}(1+\epsilon R(u)e^{-i\omega t})d\Omega^{2}_{d-2},$
where $d\Omega^{2}_{d-2}$ is the d2d-2-dimensional unit sphere and the de Sitter (cosmological) horizon is located at $u=1$. Note that after the coordinate transformation $u=\ell^{2}/r^{2}$, the metric with $\epsilon=0$ reduces to the standard static chart of $d$-dimensional de Sitter spacetime, where $u=\infty$ corresponds to the origin $r=0$. By the traceless and transverse conditions (4.6), we obtain the following three constraint equations
@@
$-(d-1)S+u(S^{\prime}-i\hat{\omega}T)=0,$ (5.2b)
$(2d-4)(1-u)^{2}R+i\hat{\omega}S+2(1-u)\{T+(u-2)U+2u(1-u)U^{\prime}\}=0,$ (5.2c)
@@
$i(\hat{m}^{2}+(d-2)\hat{\omega}^{2})S-2\hat{\omega}[(1+\hat{\omega}^{2})uT+\{2-d+(1-u)\hat{m}^{2}+(d-1+\hat{\omega}^{2})u\}U]=0.$ (5.3)
@@
$Z^{\prime\prime}+\left(\frac{1}{u-1}-\frac{d+1}{2u}\right)Z^{\prime}+\frac{1}{4u^{2}(u-1)^{2}}\left\{\hat{\omega}^{2}u-(2d+2+\hat{m}^{2})(u-1)\right\}Z=0.$ (5.4)
@@
$V(u):=\dfrac{(d+2)(d+4)}{4}\dfrac{f^{2}}{u}-(d+2)ff^{\prime}+\dfrac{f}{u}\left[2(d+1)+\hat{m}^{2}\right]\,$, (5.5b)
where f=u1f=u-1. Under the inner product defined by
@@ 5.2 Minkowski spacetime
$\Delta_{3}(x)=-\frac{\text{sgn}(t)}{2\pi}\theta(\sigma^{2})\frac{\cosh(|m|\sqrt{\sigma^{2}})}{\sqrt{\sigma^{2}}},$ (5.44a)
$\Delta_{4}(x)=-\frac{\text{sgn}(t)}{2\pi}\left(\delta(\sigma^{2})+\frac{|m^{2}|}{2}\theta(\sigma^{2})\frac{I_{1}(|m|\sqrt{\sigma^{2}})}{|m|\sqrt{\sigma^{2}}}\right),$ (5.44b)
$\Delta_{5}(x)=-\frac{\text{sgn}(t)}{(2\pi)^{2}\sigma^{2}}\Bigg[\delta(\sigma^{2})\cosh(|m|\sqrt{\sigma^{2}})$
$\hskip 85.35826pt+\theta(\sigma^{2})\left\{|m|\sinh(|m|\sqrt{\sigma^{2}})-\frac{\cosh(|m|\sqrt{\sigma^{2}})}{2\sigma^{2}}\right\}\Bigg],$ (5.44c)
@@ Appendix C The invariant delta functions $\Delta_{d}(x)$
$\Delta_{2}(x)=-\int^{\infty}_{-\infty}\frac{\sin(\omega_{q}t)}{2\pi\omega_{q}}\cos(qx)dq=-\frac{\text{sgn}(t)}{2}\theta(\sigma^{2})J_{0}(m|\sigma|),$ (C.2a)
$\Delta_{3}(x)=-\int^{\infty}_{0}\frac{q\sin(\omega_{q}t)}{2\pi\omega_{q}}J_{0}(qr)dq=-\frac{\text{sgn}(t)}{2\pi}\theta(\sigma^{2})\frac{\cos(m\sqrt{\sigma^{2}})}{\sqrt{\sigma^{2}}},$ (C.2b)
@@
$\Delta_{2}(x)=-\frac{\text{sgn}(t)}{2}\theta(\sigma^{2})I_{0}(|m||\sigma|),$ (C.4a)
$\Delta_{3}(x)=-\frac{\text{sgn}(t)}{2\pi}\theta(\sigma^{2})\frac{\cosh(|m|\sqrt{\sigma^{2}})}{\sqrt{\sigma^{2}}},$ (C.4b)
@@
$\Delta_{4}(x)=-\frac{1}{2\pi^{2}r}\int^{\infty}_{0}q\frac{\sin(\omega_{q}t)}{\omega_{q}}\sin qr\,dq=-\frac{1}{2\pi r}\frac{\partial}{\partial r}\Delta_{2}(x),$ (C.5a)
$\Delta_{5}(x)=-\frac{1}{4\pi^{2}r}\int^{\infty}_{0}q^{2}\frac{\sin(\omega_{q}t)}{\omega_{q}}J_{1}(qr)\,dq=-\frac{1}{2\pi r}\frac{\partial}{\partial r}\Delta_{3}(x),$ (C.5b)

### v2

@@ 4 The perturbed solutions with negative mass-squared
$\delta\langle~{{\cal T}_{\mu}}^{\nu}$$\rangle$ $=$ $-\frac{\epsilon L^{4}}{144\pi G_{6}\ell^{5}}\,p\,\hat{m}^{2}(\hat{m}^{2}-3){H_{\mu}}^{\nu}$ (4.20)
@@ 5.1.1 Stability analysis in the static region of de Sitter spacetime
$ds_{d}^{2}$ $=$ $-\frac{f(u)}{u}(1+\epsilon T(u)e^{-i\omega t})dt^{2}+\frac{\ell}{2u^{2}f(u)}\epsilon S(u)e^{-i\omega t}dtdu$ (5.1)
$+\frac{\ell^{2}}{4u^{2}f(u)}(1+\epsilon U(u)e^{-i\omega t})du^{2}+\frac{\ell^{2}}{u}(1+\epsilon R(u)e^{-i\omega t})d\Omega^{2}_{d-2},$
where f(u)=u1f(u)=u-1, and $d\Omega^{2}_{d-2}$ is the d2d-2-dimensional unit sphere. The de Sitter (cosmological) horizon is located at $u=1$. Note that after the coordinate transformation $u=\ell^{2}/r^{2}$, the metric with $\epsilon=0$ reduces to the standard static chart of $d$-dimensional de Sitter spacetime, where $u=\infty$ corresponds to the origin $r=0$. By the traceless and transverse conditions (4.6), we obtain the following three constraint equations
@@
$u\,S^{\prime}-\frac{d+1}{2}\,S-i\,\hat{\omega}\,u\,T=0,$ (5.2b)
$u\,U^{\prime}-\frac{1}{2}\,\left(d-1-\frac{1}{f}\right)\,U+\frac{i\,\hat{\omega}}{4\,f^{2}}\,S-\frac{1+f}{2\,f}\,T=0,$ (5.2c)
@@
$\left\{\hat{m}^{2}+(d-2)\,\hat{\omega}^{2}\right\}\,S+2\,(i\,\hat{\omega})\,\left\{(\hat{\omega}^{2}+1)\,u\,(T+U)-(\hat{m}^{2}-d+2)\,f\,U\right\}=0.$ (5.3)
@@
$Z^{\prime\prime}+\left(\frac{1}{f}-\frac{d+1}{2\,u}\right)\,Z^{\prime}+\frac{\hat{\omega}^{2}\,u-\left\{\hat{m}^{2}+2\,(d+1)\right\}\,f}{4\,u^{2}\,f^{2}}\,Z=0.$ (5.4)
@@
$V(u):=\dfrac{(d+2)(d+4)}{4}\dfrac{f^{2}}{u}-(d+2)ff^{\prime}+\dfrac{f}{u}\left[2(d+1)+\hat{m}^{2}\right]\,.$ (5.5b)
Under the inner product defined by
@@ 5.2 Minkowski spacetime
$\Delta_{3}(x)=-\frac{\text{sgn}(t)}{2\,\pi}\,\theta(\sigma^{2})\,\frac{\cosh(|m|\,\sigma)}{\sigma},$ (5.44a)
$\Delta_{4}(x)=-\frac{\text{sgn}(t)}{2\pi}\,\left(\delta(\sigma^{2})+\frac{|m^{2}|}{2}\,\theta(\sigma^{2})\,\frac{I_{1}(|m|\,\sigma)}{|m|\,\sigma}\right),$ (5.44b)
$\Delta_{5}(x)=-\frac{\text{sgn}(t)}{2\,\pi^{2}\,\sigma}\left[\,\delta(\sigma^{2})\,\cosh(|m|\,\sigma)+\frac{|m|}{2\,\sigma}\,\theta(\sigma^{2})\,\left\{\sinh(|m|\,\sigma)-\frac{\cosh(|m|\,\sigma)}{|m|\,\sigma}\right\}\,\right],$ (5.44c)
@@ Appendix C The invariant delta functions $\Delta_{d}(x)$
$\Delta_{2}(x)=-\int^{\infty}_{-\infty}\frac{dq}{2\,\pi}\,\frac{\sin(\omega_{q}t)}{\omega_{q}}\,\cos(qx)=-\frac{\text{sgn}(t)}{2}\,\theta(\sigma^{2})\,J_{0}(m\sigma),$ (C.2a)
$\Delta_{3}(x)=-\int^{\infty}_{0}\frac{dq}{2\,\pi}\,q\,\frac{\sin(\omega_{q}t)}{\omega_{q}}\,J_{0}(qr)=-\frac{\text{sgn}(t)}{2\,\pi}\,\theta(\sigma^{2})\,\frac{\cos(m\sigma)}{\sigma},$ (C.2b)
@@
$\Delta_{2}(x)=-\frac{\text{sgn}(t)}{2}\,\theta(\sigma^{2})\,I_{0}(|m|\,\sigma),$ (C.4a)
$\Delta_{3}(x)=-\frac{\text{sgn}(t)}{2\,\pi}\,\theta(\sigma^{2})\,\frac{\cosh(|m|\,\sigma)}{\sigma},$ (C.4b)
@@
$\Delta_{4}(x)=-\frac{1}{\pi\,r}\,\int^{\infty}_{0}\frac{dq}{2\,\pi}\,q\frac{\sin(\omega_{q}t)}{\omega_{q}}\,\sin(qr)=-\frac{1}{2\,\pi\,r}\,\frac{\partial}{\partial r}\,\Delta_{2}(x),$ (C.5a)
$\Delta_{5}(x)=-\frac{1}{2\,\pi\,r}\,\int^{\infty}_{0}\frac{dq}{2\,\pi}\,q^{2}\,\frac{\sin(\omega_{q}t)}{\omega_{q}}\,J_{1}(qr)=-\frac{1}{2\,\pi\,r}\,\frac{\partial}{\partial r}\,\Delta_{3}(x),$ (C.5b)